# Gravitational Wave Spectral Shapes as a probe of Long Lived Right-handed Neutrinos, Leptogenesis and Dark Matter: Global versus Local BLB-L Cosmic Strings

### v1

@@ IV.3 High-Scale Leptogenesis
To find the minimum right-handed neutrino mass for successful leptogenesis with a period of matter domination, we wish to minimise the entropy dilution factor. Using Eq. 57, this will occur when decays begin just when matter domination begins so $\rho_{R}(t_{\rm dec})\approx\rho_{N}(t_{\rm dec})$.
@@
Assuming a hierarchical mass spectrum of the right-handed neutrinos $M_{1}\ll M_{2},M_{3}$, and that the dominant lepton asymmetry arises from the decays of $N_{1}$, the CP asymmetry reduces to Fong_2012
@@
subject to our condition for matter domination Eq. 32
@@
We performed a scan of $\tilde{m}_{1}$ inputs using a maximising function and found that even with the constraint, the maximum $\epsilon$ was the Davidson-Ibarra bound Davidson:2002qv proportional to the mass of the lightest right-handed neutrino and the heaviest active neutrino,
@@
We present a benchmark in Figure 14 for successful leptogenesis in this regime, showing the evolution of the baryon asymmetry through an intermediate period of matter domination. The start and end of matter domination are marked, providing testable scales that could be probed by future gravitational-wave experiments.
Figure 14: Benchmark for successful leptogenesis showing the evolution of the baryon asymmetry together with the right-handed-neutrino and radiation energy densities through a period of matter domination. The vertical dashed lines indicate the onset and end of matter domination. Parameters: $M_{1}=10^{11}\rm\ GeV$, $\tilde{m}_{1}=10^{-11}\rm\ eV$, and $\epsilon=\epsilon^{\rm DI}/5$. Note the kink around $z=10^{2}$ is due to a change in sign.
A complete scan of the $(M_{1},\tilde{m}_{1})$ parameter space, assuming the maximum CP asymmetry and a hierarchical mass spectrum, is shown in figure 15 for both a thermal initial abundance and for a large non-thermal initial abundance with a period matter domination. The scan shows that for all initial abundances, the baryon asymmetry increases with right-handed neutrino and effective neutrino mass. This is due to $\epsilon$ scaling with $M_{1}$ and the entropy dilution scaling inversely with $\tilde{m}$. A larger initial abundance broadens the parameter space in which leptogenesis can be successful, allowing for non-thermal production to lower the bound on the right-handed neutrino mass as well as the bound on the effective neutrino mass. The black dashed line indicates the observed baryon asymmetry; leptogenesis is viable in the region of parameter space above this line.
Figure 15: Parameter scans for successful leptogenesis with a period matter domination. The left panel corresponds to a thermal initial abundance, $Y_{N}^{i}=3.9\times 10^{-3}$, while the right panel assumes a non-thermal initial abundance, $Y_{N}^{i}=1$. The black dashed line denotes the observed baryon asymmetry; parameter space above this line yields successful leptogenesis.
@@
$M_{1}\;=\;\frac{6.39\times 10^{8}}{c}\,\left(1+3.726\times 10^{-6}\,\tilde{m}^{-0.67}\right)^{3/4}\;\;\text{GeV}.$ (72)
@@ IV.4 Low-Scale Leptogenesis
where in the last step we have bounded the numerator with the Cauchy-Schwartz inequality. $Im$ denotes imaginary component. As discussed in Sec. IV.1, we remain in the non-resonant regime to avoid the need for a regulator and the associated theoretical ambiguities of the resonant case, where the CP asymmetry is highly sensitive to the mass splitting and the chosen regularisation scheme. To remain in the non-resonant regime, the mass splitting must greatly exceed both decay widths Pilaftsis_2004 ; Moffat_2018 ,
@@
The state with the smaller effective mass $\tilde{m}$ saturates the universal ceiling $|\epsilon_{i}|\lesssim 1/200$, while the other is further suppressed by the ratio of effective neutrino masses. Since the earlier–decaying right-handed neutrino’s asymmetry is erased by washout processes involving the longer–lived species, only the decay of the right-handed neutrino with the smallest effective neutrino mass is relevant. In this regime, the CP asymmetry follows the universal bound $\epsilon<1/200$. As we are dealing with small right-handed neutrino masses, we cannot assume all the right-handed neutrinos have decayed by electroweak symmetry breaking; instead, we must solve the full Boltzmann equations with this bound on $\epsilon$ numerically up to the electroweak symmetry breaking temperature $T_{\rm EW}=130\rm\ GeV$ Ulysses2 . We performed a scan and show the parameter space in figure 16. As in vanilla leptogenesis, the baryon asymmetry grows with both the right-handed-neutrino mass and the effective neutrino mass. Unlike the vanilla case, however, successful leptogenesis is realised over a much larger region of parameter space in the near-resonant case. This renders near-resonant leptogenesis far more testable than the vanilla scenario through gravitational wave observations.
@@
In our analysis, we restrict the scan to $M_{1}\gtrsim 10^{4}\,\text{GeV}$. Pushing to lower masses introduces two sources of uncertainty. First, for $M_{1}\lesssim 10^{4}\,\text{GeV}$ the onset of significant decays tends to occur only after electroweak symmetry breaking, where our treatment of rates becomes numerically unreliable. Second, in this low-mass regime, the dynamics transition away from standard vanilla leptogenesis and approach the Akhmedov-Rubakov-Smirnov (ARS) mechanism Akhmedov_1998 ; Drewes_2018 ; Klaric_2021 , where lepton asymmetry is generated by oscillations of nearly degenerate right-handed neutrinos. Since our focus is on conventional leptogenesis rather than ARS leptogenesis, we conservatively impose the cut-off at $M_{1}=10^{4}\,\text{GeV}$.
@@ IV.5 Primordial Gravitational Wave Tests of Leptogenesis
Figure 17: Detectable regions of the mass parameter space for leptogenesis with an intermediate period of matter domination. The detectable kink in the gravitational-wave background arises from cosmic strings formed by the breaking of a global (left panel) or local (right panel) $U(1)_{B-L}$ symmetry. In both cases, the outer edges of the hierarchical regime, corresponding to the Davidson Ibarra bound, are just within reach, while a substantial fraction of the near-resonant parameter space is also detectable.
Figure 17 shows that gravitational waves can probe only the outer edges of the hierarchical leptogenesis parameter space, while in the near-resonant regime they are sensitive to a much broader region of viable models. There are two interesting observables from the GWB spectral shape, $f_{\rm dom}$ and $f_{\rm brk}$. The first of these frequencies of interest, $f_{\rm dom}$, is characteristic of the onset of matter domination and is therefore determined by the right-handed neutrino mass $M_{1}$. In the hierarchical regime, large $M_{1}$ values $M_{1}>\mathcal{O}(10^{10})\rm\ GeV$ lead to very high $T_{\text{dom}}$, which shifts $f_{\text{dom}}$ beyond the reach of next-generation detectors. Consequently, only the edges of the hierarchical parameter space, with smaller $M_{1}$, produces a detectable feature. In the near-resonant regime, successful leptogenesis can occur for much smaller right-handed-neutrino masses, moving $T_{\rm dom}$ to lower values and placing $f_{\text{dom}}$ well within the sensitivity range of interferometers. The near-resonant scenario, therefore, provides a broad and accessible observational target.
The other frequency, $f_{\text{brk}}$, is characteristic of the end of matter domination at $T_{\text{end}}$, which scales with $M_{1}$ and depends on the effective neutrino mass $\tilde{m}_{1}$ through the decay rate of the right-handed neutrinos. Shorter periods of matter domination are more favourable for detection, since they produce higher break frequencies within the sensitivity range of future interferometers. Shorter periods, corresponding to $\tilde{m}_{1}$ slightly below but close to $10^{-8}\ \mathrm{eV}$, yield a narrow and well-defined feature that falls within the sensitivity bands of upcoming detectors. These frequencies, therefore, encode both the scale and the duration of sterile-neutrino domination. Importantly, such regions cannot be accessed through collider or low-energy neutrino experiments; gravitational-wave observations provide the only direct probe of this otherwise hidden sector of leptogenesis.
@@ V Gravitational Wave Tests of Dark Matter Formation
In addition to explaining neutrino masses and baryogenesis, the Type-I seesaw framework can be naturally extended to accommodate dark matter cirelli2024darkmatter . We consider such extensions in this section with an eye on how the possibility of an early period of matter domination affects the scenario. Cosmological observations firmly establish a non-baryonic component with relic density $\Omega_{\rm DM}h^{2}\simeq 0.12$ cirelli2024darkmatter . While thermal freeze-out is the canonical paradigm cirelli2024darkmatter ; Bertone_2005 ; Kolb:1990vq ; Weinberg:2008zzc ; Zeldovich:1965gev ; Bernstein:1985th ; Scherrer:1985zt ; Gondolo:1990dk ; Hisano_2007 ; Cirelli_2007 , non-thermal production via late decays provides a robust alternative that naturally arises in neutrino-mass models Falkowski_2011 ; Dutta_Banik_2021 ; Barman_2022 . In this work, the right-handed neutrino $N$ drives an early matter-dominated epoch and subsequently decays out of equilibrium into both the Standard Model and a dark sector. Concretely, we augment the BLB-L Type I seesaw Lagrangian, Eq. 12, by a Yukawa interaction
@@ V.1 Lower bound on Dark Matter Mass
If we assume that after dark matter production, there is standard cosmology, the observed dark matter yield is inversely proportional to the dark matter mass  cirelli2024darkmatter ,
@@
Here $\Omega_{\mathrm{DM}}$ is the present abundance and $M_{\rm DM}$ is the dark matter mass. The inverse scaling with $M_{\mathrm{DM}}$ implies that any dilution of the yield, in our case due to entropy injection during an early period of matter domination, must be compensated by a correspondingly larger dark matter mass in order to reproduce the observed relic abundance. If we, for now, neglect decays into the Standard Model particles, assume that all right-handed neutrinos decay into dark matter, we can obtain a lower bound on the dark matter mass under the condition of minimal entropy injection from the matter-domination era, while the upper bound comes from the necessity for decays to be allowed. For thermal production  [^5] In this section, we restrict our analysis to a thermal initial abundance of right-handed neutrinos. this lower bound is weaker than Lymann-$\alpha$ bounds on warm dark matter species Dvorkin_2021 ; Decant_2022 ; Berbig:2023yyy , the range of allowed masses of dark matter is then
@@
and matter domination condition following the logic of section II is a condition on the Yukawa coupling and the mass,
@@
These bounds are easily satisfied for even a large right-handed neutrino mass, showing that we can achieve successful dark matter and a period of matter domination for high-scale masses. We performed a systematic scan over the parameter space $(y_{\rm DM},M)$ in order to quantify the impact of the dark-sector Yukawa coupling and the right-handed neutrino mass on the duration of the matter-dominated epoch and the entropy injection. The results found that both $N_{e}$ and, as such, the entropy dilution increased with right-handed neutrino mass and decreased with Yukawa coupling; the results for $N_{e}$ are shown in the left panel of Fig. 18.
Figure 18: Left: $N_{e}$ as a function of $y_{DM}$ and $M$. Right: Dark matter mass assuming a branching ratio $\approx 1$. A longer period of matter domination increases entropy dilution, which suppresses the dark matter yield and thus requires a larger dark matter mass. The black dashed line indicates the approximate warm dark matter mass bound. For the assumed branching ratio, the viability of the dark matter candidate requires a period of matter domination. In both plots, only parameter points that feature a period of matter domination are shown.
By matching the results to the observed dark matter abundance, we determine the parametric dependence of the dark matter mass on the right-handed neutrino parameters. As shown in the right panel of Fig. 18, the mass of the dark matter candidate grows with increasing right-handed neutrino mass and decreases with larger Yukawas. This behaviour reflects the fact that a longer period of matter domination enhances entropy injection, thereby reducing the dark matter yield and necessitating a heavier dark matter mass. The plot assumes a branching ratio of unity. In this case, a viable dark matter candidate only emerges if the universe undergoes a period of matter domination and the subsequent entropy dilution, this is shown in the $M_{\rm DM}>\mathcal{O}(10)$ keV bound in the right panel of Fig  18. More generally, if the branching ratio into dark matter is below $1\%$, this requirement is lifted; but once it exceeds $1\%$, early matter domination and the subsequent entropy dilution become essential for dark matter to remain viable.
@@ V.2 Lower bound on Asymmetric Dark Matter Mass
One of the most striking features of the Universe is the near coincidence between the baryonic and dark matter energy densities, $\Omega_{\mathrm{DM}}/\Omega_{B}\simeq 5$. Despite their apparently distinct origins, this numerical proximity hints at a common mechanism that simultaneously generates both components of cosmic matter. Asymmetric dark matter (ADM) models Falkowski_2011 ; Dutta_Banik_2021 ; Barman_2022 offer a natural explanation: the dark matter relic abundance arises from an initial matter–antimatter asymmetry rather than thermal freeze-out, mirroring the generation of the baryon asymmetry through leptogenesis. In such frameworks, the CP-violating decays of heavy right-handed neutrinos directly produce lepton asymmetries in both the visible and dark sectors. The lepton asymmetry in the Standard Model is subsequently converted into baryon number via sphaleron processes, while the corresponding dark asymmetry determines the relic density of the dark sector. This mechanism elegantly unifies the explanations for neutrino masses, the baryon asymmetry of the Universe, and the dark matter abundance within a single extension of the Standard Model, suggesting that the cosmic composition of matter ultimately traces back to the dynamics of sterile neutrino decays.
@@
Following Falkowski_2011 , the dark matter abundance will depend on the dark matter CP asymmetry, efficiency, as well as initial abundance and entropy dilution
@@
For hierarchical right-handed neutrino masses, the parameter $\epsilon_{\chi}$ is dependent on these masses and the Yukawa couplings Falkowski_2011 , and can be expressed as,
@@
which imposing the condition $M_{\rm DM}<M_{1}$ has the bound for successful matter domination and leptogenesis and asymmetric dark matter to be, $\mathcal{Y}_{2}\gtrsim 10^{-14}$ showing these conditions are trivially satisfied.

### v2

@@ IV.3 High-Scale Leptogenesis
To find the minimum right-handed neutrino mass for successful leptogenesis with a period of early matter domination, we wish to minimise the entropy dilution factor. Using Eq. 63, this will occur when decays begin just when early matter domination begins so $\rho_{R}(t_{\rm dec})\approx\rho_{N}(t_{\rm dec})$.
@@
Assuming a hierarchical mass spectrum of the right-handed neutrinos $M_{1}\ll M_{2},M_{3}$, and that the dominant lepton asymmetry arises from the decays of $N_{1}$, the CP asymmetry reduces to Fong et al. (2012)
@@
subject to our condition for early matter domination Eq. 32
@@
We performed a scan of $\tilde{m}_{1}$ inputs using a maximising function and found that even with the constraint, the maximum $\epsilon$ was the Davidson-Ibarra bound Davidson and Ibarra (2002) proportional to the mass of the lightest right-handed neutrino and the heaviest active neutrino,
@@
We present a benchmark in Figure 14 for successful leptogenesis in this regime, showing the evolution of the baryon asymmetry through an intermediate period of early matter domination. The start and end of early matter domination are marked, providing testable scales that could be probed by future gravitational-wave experiments.
Figure 14: Benchmark for successful leptogenesis showing the evolution of the baryon asymmetry together with the right-handed-neutrino and radiation energy densities through a period of early matter domination. The vertical dashed lines indicate the onset and end of early matter domination. Parameters: $M_{1}=10^{11}\rm\ GeV$, $\tilde{m}_{1}=10^{-11}\rm\ eV$, and $\epsilon=\epsilon^{\rm DI}/5$. Note the kink around $z=10^{2}$ is due to a change in sign.
A complete scan of the $(M_{1},\tilde{m}_{1})$ parameter space, assuming the maximum CP asymmetry and a hierarchical mass spectrum, is shown in figure 15 for both a thermal initial abundance and for a large non-thermal initial abundance with a period early matter domination. The scan shows that for all initial abundances, the baryon asymmetry increases with right-handed neutrino and effective neutrino mass. This is due to $\epsilon$ scaling with $M_{1}$ and the entropy dilution scaling inversely with $\tilde{m}$. A larger initial abundance broadens the parameter space in which leptogenesis can be successful, allowing for non-thermal production to lower the bound on the right-handed neutrino mass as well as the bound on the effective neutrino mass. The black dashed line indicates the observed baryon asymmetry; leptogenesis is viable in the region of parameter space above this line.
Figure 15: Parameter scans for successful leptogenesis with a period of early matter domination. The left panel corresponds to a thermal initial abundance, $Y_{N}^{i}=3.9\times 10^{-3}$, while the right panel assumes a non-thermal initial abundance, $Y_{N}^{i}=1$. The black dashed line denotes the observed baryon asymmetry; parameter space above this line yields successful leptogenesis.
@@
$M_{1}\;=\;\frac{6.39\times 10^{8}}{c}\,\left(1+3.726\times 10^{-6}\,\left(\frac{\tilde{m}}{\rm eV}\right)^{-0.67}\right)^{3/4}\;\;\text{GeV}.$ (78)
@@ IV.4 Low-Scale Leptogenesis
where in the last step we have bounded the numerator with the Cauchy-Schwartz inequality. $Im$ denotes imaginary component. As discussed in Sec. IV.1, we remain in the non-resonant regime to avoid the need for a regulator and the associated theoretical ambiguities of the resonant case, where the CP asymmetry is highly sensitive to the mass splitting and the chosen regularisation scheme. To remain in the non-resonant regime, the mass splitting must greatly exceed both decay widths Pilaftsis and Underwood (2004); Moffat et al. (2018),
@@
The state with the smaller effective mass $\tilde{m}$ saturates the universal ceiling $|\epsilon_{i}|\lesssim 1/200$, while the other is further suppressed by the ratio of effective neutrino masses. Since the earlier–decaying right-handed neutrino’s asymmetry is erased by washout processes involving the longer–lived species, only the decay of the right-handed neutrino with the smallest effective neutrino mass is relevant. In this regime, the CP asymmetry follows the universal bound $\epsilon<1/200$. As we are dealing with small right-handed neutrino masses, we cannot assume all the right-handed neutrinos have decayed by electroweak symmetry breaking; instead, we must solve the full Boltzmann equations with this bound on $\epsilon$ numerically up to the electroweak symmetry breaking temperature $T_{\rm EW}=130\rm\ GeV$ Granelli et al. (2023). We performed a scan and show the parameter space in figure 16. As in vanilla leptogenesis, the baryon asymmetry grows with both the right-handed-neutrino mass and the effective neutrino mass. Unlike the vanilla case, however, successful leptogenesis is realised over a much larger region of parameter space in the near-resonant case. This renders near-resonant leptogenesis far more testable than the vanilla scenario through gravitational wave observations.
@@
In our analysis, we restrict the scan to $M_{1}\gtrsim 10^{4}\,\text{GeV}$. This is because in this low-mass regime, the dynamics transition away from standard vanilla leptogenesis and approach the Akhmedov-Rubakov-Smirnov (ARS) mechanism Akhmedov et al. (1998); Drewes et al. (2018); Klaric et al. (2021), where lepton asymmetry is generated by oscillations of nearly degenerate right-handed neutrinos. Since our focus is on conventional leptogenesis rather than ARS leptogenesis, we conservatively impose the cut-off at $M_{1}=10^{4}\,\text{GeV}$.
@@ IV.5 Primordial Gravitational Wave Tests of Leptogenesis
Figure 17: Detectable regions of the mass parameter space for leptogenesis with an intermediate period of early matter domination. The detectable kink in the gravitational-wave background arises from cosmic strings formed by the breaking of a global (left panel) or local (right panel) $U(1)_{B-L}$ symmetry. In both cases, the outer edges of the hierarchical regime, corresponding to the Davidson Ibarra bound, are just within reach, while a substantial fraction of the near-resonant parameter space is also detectable.
Figure 17 shows that gravitational waves can probe only the outer edges of the hierarchical leptogenesis parameter space, while in the near-resonant regime they are sensitive to a much broader region of viable models. There are two interesting observables from the GWB spectral shape, $f_{\rm dom}$ and $f_{\rm brk}$. The first of these frequencies of interest, $f_{\rm dom}$, is characteristic of the onset of early matter domination and is therefore determined by the right-handed neutrino mass $M_{1}$. In the hierarchical regime, large $M_{1}$ values $M_{1}>\mathcal{O}(10^{10})\rm\ GeV$ lead to very high $T_{\text{dom}}$, which shifts $f_{\text{dom}}$ beyond the reach of next-generation detectors. Consequently, only the edges of the hierarchical parameter space, with smaller $M_{1}$, produces a detectable feature. In the near-resonant regime, successful leptogenesis can occur for much smaller right-handed-neutrino masses, moving $T_{\rm dom}$ to lower values and placing $f_{\text{dom}}$ well within the sensitivity range of interferometers. The near-resonant scenario, therefore, provides a broad and accessible observational target.
The other frequency, $f_{\text{brk}}$, is characteristic of the end of early matter domination at $T_{\text{end}}$, which scales with $M_{1}$ and depends on the effective neutrino mass $\tilde{m}_{1}$ through the decay rate of the right-handed neutrinos. Shorter periods of early matter domination are more favourable for detection, since they produce higher break frequencies within the sensitivity range of future interferometers. Shorter periods, corresponding to $\tilde{m}_{1}$ slightly below but close to $10^{-8}\ \mathrm{eV}$, yield a narrow and well-defined feature that falls within the sensitivity bands of upcoming detectors. These frequencies, therefore, encode both the scale and the duration of sterile-neutrino domination. Importantly, such regions cannot be accessed through collider or low-energy neutrino experiments; gravitational-wave observations provide the only direct probe of this otherwise hidden sector of leptogenesis.
@@ V Gravitational Wave Tests of Dark Matter Formation
In addition to explaining neutrino masses and baryogenesis, the Type-I seesaw framework can be naturally extended to accommodate dark matter Cirelli et al. (2024). We consider such extensions in this section with an eye on how the possibility of an early period of matter domination affects the scenario. Cosmological observations firmly establish a non-baryonic component with relic density $\Omega_{\rm DM}h^{2}\simeq 0.12$ Cirelli et al. (2024). While thermal freeze-out is the canonical paradigm Cirelli et al. (2024); Bertone et al. (2005); Kolb and Turner (2019); Weinberg (2008); Zeldovich (1965); Bernstein et al. (1985); Scherrer and Turner (1986); Gondolo and Gelmini (1991); Hisano et al. (2007); Cirelli et al. (2007), non-thermal production via late decays provides a robust alternative that naturally arises in neutrino-mass models Falkowski et al. (2011); Dutta Banik et al. (2021); Barman et al. (2022). In this work, the right-handed neutrino $N$ drives an early matter-dominated epoch and subsequently decays out of equilibrium into both the Standard Model and a dark sector. Concretely, we augment the BLB-L Type I seesaw Lagrangian, Eq. 12, by a Yukawa interaction
@@ V.1 Lower bound on Dark Matter Mass
If we assume that after dark matter production, there is standard cosmology, the observed dark matter yield is inversely proportional to the dark matter mass  Cirelli et al. (2024),
@@
Here $\Omega_{\mathrm{DM}}$ is the present abundance and $M_{\rm DM}$ is the dark matter mass. The inverse scaling with $M_{\mathrm{DM}}$ implies that any dilution of the yield, in our case due to entropy injection during an early period of matter domination, must be compensated by a correspondingly larger dark matter mass in order to reproduce the observed relic abundance. If we, for now, neglect decays into the Standard Model particles, assume that all right-handed neutrinos decay into dark matter, we can obtain a lower bound on the dark matter mass under the condition of minimal entropy injection from the matter-domination era, while the upper bound comes from the necessity for decays to be allowed. For thermal production  [^7] In this section, we restrict our analysis to a thermal initial abundance of right-handed neutrinos. this lower bound is weaker than Lyman-$\alpha$ bounds on warm dark matter species Dvorkin et al. (2021); Decant et al. (2022); Berbig and Ghoshal (2023), the range of allowed masses of dark matter is then
@@
and early matter domination condition following the logic of section II is a condition on the Yukawa coupling and the mass,
@@
These bounds are easily satisfied for even a large right-handed neutrino mass, showing that we can achieve successful dark matter and a period of early matter domination for high-scale masses. We performed a systematic scan over the parameter space $(y_{\rm DM},M)$ in order to quantify the impact of the dark-sector Yukawa coupling and the right-handed neutrino mass on the duration of the matter-dominated epoch and the entropy injection. The results found that both $N_{e}$ and, as such, the entropy dilution increased with right-handed neutrino mass and decreased with Yukawa coupling; the results for $N_{e}$ are shown in the left panel of Fig. 18.
Figure 18: Left: $N_{e}$ as a function of $y_{DM}$ and $M$. Right: Dark matter mass assuming a branching ratio $\approx 1$. A longer period of early matter domination increases entropy dilution, which suppresses the dark matter yield and thus requires a larger dark matter mass. The black dashed line indicates the approximate warm dark matter mass bound. For the assumed branching ratio, the viability of the dark matter candidate requires a period of early matter domination. In both plots, only parameter points that feature a period of early matter domination are shown.
By matching the results to the observed dark matter abundance, we determine the parametric dependence of the dark matter mass on the right-handed neutrino parameters. As shown in the right panel of Fig. 18, the mass of the dark matter candidate grows with increasing right-handed neutrino mass and decreases with larger Yukawas. This behaviour reflects the fact that a longer period of early matter domination enhances entropy injection, thereby reducing the dark matter yield and necessitating a heavier dark matter mass. The plot assumes a branching ratio of unity. In this case, a viable dark matter candidate only emerges if the universe undergoes a period of early matter domination and the subsequent entropy dilution, this is shown in the $M_{\rm DM}>\mathcal{O}(10)$ keV bound in the right panel of Fig  18. More generally, if the branching ratio into dark matter is below $1\%$, this requirement is lifted; but once it exceeds $1\%$, early matter domination and the subsequent entropy dilution become essential for dark matter to remain viable.
@@ V.2 Lower bound on Asymmetric Dark Matter Mass
One of the most striking features of the Universe is the near coincidence between the baryonic and dark matter energy densities, $\Omega_{\mathrm{DM}}/\Omega_{B}\simeq 5$. Despite their apparently distinct origins, this numerical proximity hints at a common mechanism that simultaneously generates both components of cosmic matter. Asymmetric dark matter (ADM) models Falkowski et al. (2011); Dutta Banik et al. (2021); Barman et al. (2022) offer a natural explanation: the dark matter relic abundance arises from an initial matter–antimatter asymmetry rather than thermal freeze-out, mirroring the generation of the baryon asymmetry through leptogenesis. In such frameworks, the CP-violating decays of heavy right-handed neutrinos directly produce lepton asymmetries in both the visible and dark sectors. The lepton asymmetry in the Standard Model is subsequently converted into baryon number via sphaleron processes, while the corresponding dark asymmetry determines the relic density of the dark sector. This mechanism elegantly unifies the explanations for neutrino masses, the baryon asymmetry of the Universe, and the dark matter abundance within a single extension of the Standard Model, suggesting that the cosmic composition of matter ultimately traces back to the dynamics of sterile neutrino decays.
@@
Following Falkowski et al. (2011), the dark matter abundance will depend on the dark matter CP asymmetry, efficiency, as well as initial abundance and entropy dilution
@@
For hierarchical right-handed neutrino masses, the parameter $\epsilon_{\chi}$ is dependent on these masses and the Yukawa couplings Falkowski et al. (2011), and can be expressed as,
@@
which imposing the condition $M_{\rm DM}<M_{1}$ has the bound for successful early matter domination and leptogenesis and asymmetric dark matter to be, $\mathcal{Y}_{2}\gtrsim 10^{-14}$ showing these conditions are trivially satisfied.